### v5

@@ Abstract.
We extend this algorithm and tool to support quotient type equivalences in Coq. Notably, since Coq lacks quotient types entirely, our extensions use Coq’s setoid machinery to represent quotients externally. Specifically, (1) our extension to the algorithm supports new changes corresponding to setoids, and (2) our extension to the tool supports this new class of changes and further automates away some of the new proof obligations. We ground our setoid extensions by way of a discussion of a corresponding manual proof repair approach in Cubical Agda, which supports quotient types and allows for some internalization of the correctness criteria for proof repair. We demonstrate our extensions on proof repair case studies for previously unsupported changes.
@@ 1. Introduction
Proof repair (Ringer, 2021) aims to simplify this process by introducing algorithms and tools that fix formal proofs in response to breaking changes. Prior work introduced a Coq plugin called Pumpkin Pi for proof repair across changes in datatypes that can be described by type equivalences (Ringer et al., 2021). Pumpkin Pi could handle only a limited class of changes corresponding to non-equivalent types by way of expressing those changes as equivalences between sigma types. The power of this was limited, however, and could not represent fundamental changes in behavior in a practical way.
In this work, we extend proof repair to support equivalences between quotient types (Section 2). Recent work in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between quotient types (Angiuli et al., 2021). However, unlike Cubical Agda, Coq lacks quotient types entirely, so one cannot use the original Pumpkin Pi transformation as-is to support this class of changes. To handle this, we represent quotient types externally using Coq’s setoid machinery, and we represent quotient type equivalences as setoid equivalences (Section 3). We then extend the proof transformation to support the newly generated equality proof obligations (Section 4).
We implement the extended algorithm by way of an extension to the implementation of Pumpkin Pi (Section 5). Our implementation includes new automation, both to support repair across this class of changes, and to automate away some of the newly generated proof obligations corresponding to this class of changes. We demonstrate our extended implementation on three case studies that cannot be handled by prior proof repair work—two that are mathematical in nature and one that deals with changes in behavior (Section 6). We ground the meaning of our extended transformation by way of a discussion comparing to a manual approach in Cubical Agda (Section 7). Our contributions are:
@@
    a discussion grounding the meaning of the above by comparing to a manual approach in Cubical Agda.
Our code is available in supplementary material.
2. Problem Definition
The problem that we explore is an extension of the proof repair problem from Pumpkin Pi. Given proofs defined over some old type $A$, and an equivalence between $A$ and some new type $B$, Pumpkin Pi repairs proofs that refer to $A$ to instead refer to $B$. In this way, it operates a lot like an automated transfer or transport method (see Section 8), but it also goes one step further: since $A$ represents an old version of the type, and $B$ a new version, repair further removes all references to $A$ in the repaired functions and proofs, so that they refer only to $B$. In contrast, transport and transfer methods typically produce terms that refer to both $A$ and $B$, for example by applying functions mapping between them.
We expand this problem to allow for the set of pairs of equivalent types $A$ and $B$ that we can repair proofs over to include a useful class of changes in types not directly describable in Coq—those that can be described by quotient type equivalences (Section 2.1). For simplicity, we start by considering this problem in a fragment of cubical type theory which supports quotient types directly. In such a context, we can internally and formally state what it means for any given repaired proof defined over $B$ to correctly correspond to the original proof over $A$—in terms of dependent path equality (Section 2.2). We will later explain how we can approximate this externally in Coq using setoids (Section 3), and we will return to cubical for a discussion at the end (Section 7).
2.1. Scope: Quotient Type Equivalences
In this paper, we perform proof repair across quotient type equivalences. We will describe what makes these quotient type equivalences soon. But in general, a type equivalence between two types $A$ and $B$ is an isomorphism (a pair of functions that are mutual inverses) satisfying a particular coherence property (Univalent Foundations Program, 2013). It is possible to form an equivalence from any isomorphism.
Like Pumpkin Pi in Coq, we consider changes in datatypes that can be described by these type equivalences, but we also consider a broader class of changes than type equivalences encode in Coq. The trick is to encode changes in datatypes as equivalences between quotient types—types equipped with an equivalence relation describing what makes two elements of that type “the same.” Since Coq cannot directly express these types, we will later approximate them externally.
For now, let us work in cubical, which has internal quotient types. We denote quotient types as $A/R$ for a type $A$ and a relation $R$. An element of this quotient is the equivalence class of any $a$ of type $A$. Given elements $[a_{1}]$ and $[a_{2}]$, we have that $[a_{1}]\equiv[a_{2}]$ whenever $R\ a_{1}\ a_{2}$, even when it is false that $a_{1}\equiv a_{2}$. For example, consider the natural numbers mod 2, $\mathbb{N}/2$. Two naturals are in the same equivalence class if they have the same parity. This type has two equivalence classes, and so every element of the type is equal to either [ 0 ] or [ 1 ]. For another element, like [ 2 ], we have a proof of equality between [ 0 ] and [ 2 ], since both are even.
Given two isomorphic types, at least one of which is a quotient type, we can form type equivalences between them in cubical just as we normally could with other isomorphic types. For example, $\mathbb{N}/2$ is isomorphic to the booleans via the map [ 0 ] $\mapsto$ false, [ 1 ] $\mapsto$ true, and this can be used to construct a type equivalence. In this paper, we further restrict ourselves to equivalences between set quotients, which have the additional constraint that all proofs of equality between elements of the quotient must themselves be equal to each other. In homotopy type theory parlance, $A/R$ must be an h-set; in broader dependent type theory parlance, uniqueness of identity proofs must provably hold for $A/R$.
Quotient type equivalences are useful for mathematics as well as for describing changes in implementation of datatypes. For example, in Section 6.2 we will show how to use quotient type equivalences to repair functions and proofs between two non-isomorphic implementations of a queue: one backed by a single list, and one backed by a pair of lists. For the latter, we will also implement an efficient dequeueing function that operates by removing an element from the front of the second list of the pair, and an efficient enqueueing function that works by appending to the front of the first list in the pair. We will then use a straightforward methodology to obtain repaired proofs over the efficient functions as well.
Supporting quotient type equivalences makes repair across this change possible for the first time. In Pumpkin Pi, some changes more interesting than vanilla type equivalences could be expressed as equivalences between sigma types; we could try that approach for quotients, but it would come with severe disadvantages. For example, we could represent queues backed by pairs of lists as the sigma type:
$\sum_{x\>:\>\text{List }A\>\times\>\text{List }A}isCanonical\ x$
where $isCanonical\ x$ is a proposition stating that $x$ is the canonical representative of its equivalence class. This type is equivalent to one list queues, but has only one representative of each equivalence class. This makes an efficient implementation of both enqueue and dequeue impossible. Quotient types empower us to obtain repaired functions and proofs that operate over the efficient implementations in the end.
2.2. Goal: Dependent Path Equality
In cubical type theory, we can in general state what it means for a repaired proof to be correctly related to the original proof. This is because cubical type theory is univalent (Cohen et al., 2018; Angiuli et al., 2017)—equivalence is equivalent to equality (Univalent Foundations Program, 2013). [^1] In cubical type theory, univalence is not an axiom, but rather follows computationally from more primitive constructs. We take advantage not of univalence directly, but of these more primitive constructive constructs.
We define correctness in terms of a generalization of path equality, the primitive notion of propositional equality in cubical. If two terms a and b of the same type A are path equal, we can produce a term of type a $\equiv$ b. Using univalence, we can obtain an isomorphism from any path equality. For example, from our isomorphism between $\mathbb{N}/2$ and the booleans, we can obtain a path equality of type $\mathbb{N}/2$equiv bool.
Path equality is actually an instantiation of the more powerful heterogeneous notion of dependent path equality. In this paper, we use Cubical Agda syntax to denote dependent path equality. Terms a : A and b : B are dependently path equal if there is a term of type PathP p a b where p is a path between A and B. For example, let p be our proof that $\mathbb{N}/2$equiv bool. Then there is a term of type PathP p [ 0 ] false because [ 0 ] maps to false under the isomorphism we used to create p. Non-dependent path equality $\equiv$ is dependent path equality at a single type, where p is the identity path.
We can state correctness of a repaired proof relative to the corresponding original version by defining the correct PathP type—the one between the original term and the repaired term at the path between their types. For example, the type PathP p [ 0 ] false can be viewed as stating that it is correct to repair [ 0 ] to false across $\mathbb{N}/2\equiv$ bool.
We will describe how to use this to define correctness more generally in Section 7. We will also formalize and prove some of these correctness theorems for the very first time, for manual proof repair in of Cubical Agda. Consistently with prior work, though, our extension to Pumpkin Pi for Coq will simply aspire to satisfy the correctness criteria from Pumpkin Pi, which are stated in an unspecified univalent metatheory, without any correctness proofs.
We will further require that repaired functions and proofs no longer contain references to the old version of those functions and proofs, defined over the old version of the type. Also consistently with prior work, we will not state this formally. But this is the primary way in which proof repair diverges from (dependent) transport: using transport to move proofs across equalities forces both equivalent types to eternally remain in the codebase. Repair lets us move on and remove the old code.
3. Trick: Setoid Equivalences
We now return from cubical to Coq, the proof assistant that our extension to Pumpkin Pi targets. In contrast with cubical, Coq’s type theory does not support quotient types at all. So how can we support proof repair across quotient type equivalences when quotient types do not even exist?
Our answer is to work with setoids—an external notion of quotients. Setoids in Coq are types paired with an equivalence relation representing equality (Sozeau, 2023). For example, we can represent our $\mathbb{N}/2$ type from Section 2.1 in Coq as the setoid (nat, mod_two), where mod_two is again equivalence modulo $2$. In this case, we call nat the carrier of the setoid. We define functions and theorems on elements of the carrier.
Coq has a Setoid type class, and any setoid defines an instance of this type class. However, our extension will not use this type class. Instead, we will understand metatheoretically that any pair of a type and an equivalence relation on that type forms a setoid, and we will use the machinery that derives from instances of the Equivalence and Proper type classes. We can view any type as a setoid this way by letting the equivalence relation be propositional equality, but more interesting setoids will use equivalence relations that do not correspond to propositional equality in Coq.
When comparing elements of a setoid, the equivalence relation is used in place of equality. As a consequence of this, users of setoids need to juggle multiple notions of equality, unlike with native quotient types where the same equality is used universally. For example, in cubical we can compare equality of elements of $\mathbb{N}/2$ using the native equality $a\equiv b$, but in Coq, when considering (nat, mod_two) as a setoid, it is too strong to claim that a = b. Instead, we compare equality of elements of the setoid (nat, mod_two) by using the equivalence relation mod_two a b.
Coq’s setoids do not enforce that functions defined on them respect the equivalence relation until a user needs to rewrite under that function in a proof, unlike with native quotient types, where well-definedness is checked statically upon writing the function. For example, we can define:
f (x : nat) : bool := eqb x 3
as a valid function for our (nat, mod_two) setoid in Coq. However, unlike with propositional equality, it is not the case that f x1 = f x2 whenever mod_two x1 x2, and we cannot rewrite f x1 to f x2 using a proof that mod_two x1 x2.
In other words, this particular choice of f is not proper. When it is the case that, for setoids (A, eqA) and (B, eqB) and a function f : A -> B, the theorem:
$\forall$ (a1 a2 : A), (eqA a1 a2) -> eqB (f a1) (f a2)
holds, we say that f is proper. A function satisfying this property defines an instance of the Proper type class in Coq. For example, consider isEven : nat -> bool, which sends even numbers to true and odd numbers to false. This is proper considering the domain as the setoid (nat, mod_two) and the codomain as the setoid (bool, =).
@@ 3. Trick: Setoid Equivalences
To represent quotient type equivalences in Coq, we define a notion of a setoid equivalence. Two setoids (A, eqA) and (B, eqB) are equivalent if there is a pair of functions f : A -> B and g : B -> A satisfying the following properties:
@@
As an example, the isEven function is one half of an equivalence between (nat, mod_two) and (bool, =). It has as an inverse that sends true to 0 and false to 1. It is easy to verify that this pair of functions satisfies the above properties.
@@ 4. Approach: Proof Term Transformation
The way that Pumpkin Pi repairs proofs across type equivalences is by directly transforming proof terms across those equivalences (Section 4.1). We extend Pumpkin Pi’s transformation to support setoid equivalences (Section 4.2).
4.1. Pumpkin Pi’s Transformation
{grammar}
[…]
¡t¿ ::= ¡v¿   —   ¡s¿   —   $\Pi$ (¡v¿ : ¡t¿) . ¡t¿   —   $\lambda$ (¡v¿ : ¡t¿) . ¡t¿   —   ¡t¿ ¡t¿   —   Ind (¡v¿ : ¡t¿){¡t¿,…,¡t¿}   —   Constr (¡i¿, ¡t¿)   —   Elim(¡t¿, ¡t¿){¡t¿,…,¡t¿}
@@ 4.1. Pumpkin Pi’s Transformation
Figure 3. Transformation for repair across $A\simeq B$ with configuration ((DepConstr, DepElim), (Eta, Iota)), from previous work (Ringer, 2021). Our work adapts and extends this transformation.
(DepConstr, DepElim) , ($\iota$, $\eta$)
for types on both sides of the equivalence. [^2] For the purposes of this paper, $\eta$, dealing with $\eta$-expansions of constructors applied to eliminators, will always be trivial, and thus we will ignore it. DepConstr and DepElim are, respectively, constructors and eliminators for each type. The constructors must generate the elements of the type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators must take the same shape, even if A and B themselves have different shapes.
To use an example from the Pumpkin Pi paper, for the type of unary naturals in Figure 2, we could choose depConstr for nat to be O and S, the constructors for nat. However, to repair to the binary naturals, we need to provide dependent constructors N. These constructors must correspond across the equivalence, so we arrive at constructors with these types:
Definition depConstrNZero : N.
Definition depConstrNSuc : N -> N.
These dependent constructors do not share the type signatures of N’s constructors, but rather are user-defined functions corresponding to nat’s constructors. Likewise, the dependent eliminator for N takes the shape of nat’s eliminator:
Definition depElimN : $\forall$ (P : N -> Type),
(P depConstrNZero) ->
($\forall$ n : N, P n -> P (depConstrNSuc n)) ->
$\forall$ n : N, P n.
The fact that both dependent eliminators must have the same shape even when the underlying types do not is exactly why we need the remaining element of the configuration: $\iota$. This gives the $\iota$-reduction rules, which specify how to reduce an application of a dependent eliminator to a dependent constructor. When the shape is of the underlying type is the same as the shape of the configuration components, as is true for for nat, this will be definitional—the proof assistant will handle it automatically. But if the inductive structure has changed, as it has N, this will be a propositional equality.
Once we have defined the components of the configuration, we are ready to do repair. First, the functions we wish to repair are converted to explicitly refer to the configuration terms. Sometimes this happens via manual user annotation, and sometimes this happens via custom unification machinery inside of Pumpkin Pi that does this automatically for some classes of changes (see Section 5.2). Then, we follow the syntactic transformation outlined in Figure 3. We will discuss what it means for this to be correct in Section 7.2.

### v6

@@ Abstract.
We extend this algorithm and tool to support quotient type equivalences in Coq. Notably, since Coq lacks quotient types entirely, our extensions use Coq’s setoid machinery to represent quotients externally. Specifically, (1) our extension to the algorithm supports new changes corresponding to setoids, and (2) our extension to the tool supports this new class of changes and further automates away some of the new proof obligations. We demonstrate our extensions on proof repair case studies for previously unsupported changes. We also perform manual proof repair in Cubical Agda, a language with a univalent metatheory, which allows us to construct the first ever internal proofs of correctness for proof repair.
@@ 1. Introduction
Proof repair (Ringer, 2021) aims to simplify this process by introducing algorithms and tools that fix formal proofs in response to breaking changes. Given an existing type $A$ and some set of functions and theorems on that type, as well a new type $B$, proof repair seeks to generate the equivalent functions and theorems defined on the type $B$. Proof repair further requires that the new functions and proofs on $B$ make no reference to the original functions and proofs on $A$, which distinguishes it from the more general proof transfer.
Prior work introduced Pumpkin Pi, a Coq plugin for proof repair across changes in datatypes that can be described by type equivalences (Ringer et al., 2021). In this work, we extend proof repair to support equivalences between quotient types (Section 2). Recent work in Cubical Agda showed that certain relations describing changes in behavior can be adjusted to equivalences between quotient types (Angiuli et al., 2021). One specific example presented in that work uses two of the most basic queue representations. The first, one list queues, enqueue elements to the front of the list and dequeue elements from the back of the list. The second, two list queues, use a pair of lists. Elements are enqueued onto the front of the first list. When dequeuing, if the second list is empty, the first list is reversed onto the second, and then in any case the front element of the second list is removed. These types are not equivalent in any natural way, because multiple two list queues correspond to a single one list queue: for instance, the one list queue l corresponds to both (l, []) and ([], rev l). Angiuli et al. quotient two list queues so that, in the quotient type, (l, []) and ([], rev l) are in the same equivalence class, which makes the types equivalent. They are then able to perform proof transfer across that equivalence.
We wish to use this same approach to implement proof repair instead of proof transfer. Unlike Cubical Agda, Coq lacks quotient types entirely, so one cannot use the original Pumpkin Pi transformation as-is to support this class of changes. To handle this, we represent quotient types externally using Coq’s setoid machinery, and we represent quotient type equivalences as setoid equivalences (Section 2). We then extend the proof transformation to support the newly generated equality proof obligations (Section 3).
By extending Pumpkin Pi to quotient equivalences, users gain both expressive power and efficiency. Prior to this work, Pumpkin Pi could capture a similar approach using sigma types instead of setoids. To do this, one would choose a canonical element for each equivalence class in the quotient, and then use the subtype of these elements instead of a quotient or setoid. However, that approach has severe drawbacks. In the two list queue example above, when using both quotient and setoids representations, the multiple representatives of each class allow for an amortized constant time dequeue operation. The subtype representation instead only allows for a linear time dequeue operation.
We implement the extended algorithm by way of an extension to the implementation of Pumpkin Pi (Section 4). Our implementation includes new automation, both to support repair across this class of changes and to automate away some of the newly generated proof obligations corresponding to this class of changes. We demonstrate our extended implementation on three case studies that cannot be handled by prior proof repair work—two that are mathematical in nature and one that deals with changes in behavior (Section 5). Finally, we define correctness for proof repair and construct internal proofs of correct repair in Cubical Agda (Section 6). Our contributions are:
@@
    the first construction of internal proofs of correct repair.
Our code is available in supplementary material.  [^1] https://github.com/uwplse/pumpkin-pi/tree/quotient-type-equivalences
2. Quotients and Setoids
Our goal is to extend Pumpkin Pi to perform repair between types equivalent up to quotients. To begin this task, we define quotient types:
Definition 2.1.
A quotient type A / eqA is defined by a type A and an equivalence relation eqA : A $\to$ A $\to$ Prop. A / eqA has one constructor, [ $\cdot$ ] : A $\to$ A / eqA, and its elements are called the equivalence classes of the elements of A. A / eqA has the additional property that, for a, b : A, [ a ] = [ b ] if eqA a b is inhabited. Eliminating an element q : A / eqA yields the underlying a : A from which it was constructed. However, the user must prove that the result of the computation is equal for any a, b : A with [ a ] = [ b ].
To give a concrete example, consider the quotient of $\mathbb{N}$ by the equivalence relation $\mathbb{N}$, where $\mathbb{N}$ n m is inhabited if n and m have the same parity. The resulting type, denoted by mathematicians as $\mathbb{N}/2$, has two elements, [ 0 ] and [ 1 ]; any other application of [ $\cdot$ ] produces a term equal to one of these. For $\mathbb{N}/2$, eliminating [ 0 ] and eliminating [ 2 ] must provably result in equal terms.
Quotient types can be challenging to implement within a language because of the equality property mentioned above. While some languages have native quotient types, many type systems, including Coq, do not support this property without the use of axioms. To circumvent this issue without introducing any axioms, we use setoids.
Definition 2.2.
A setoid is the pair (A, eqA). When the equivalence relation eqA is obvious from context, we may also call A a setoid as shorthand.
Notice that setoids have no special constructors, eliminators, or equality properties. An element of A is said to be an element of the setoid, and users of the setoid should compare elements using eqA instead of native equality. Because of this, rewriting is more difficult in setoids than in quotient types. Rewriting by an equality in a quotient type can be done using the equality eliminator, but rewriting cannot generally be done for arbitrary equivalence relations. To compensate for this, we define the notion of a proper function.
Definition 2.3.
For setoids (A, eqA) and (B, eqB), a function f : A $\to$ B is proper if, for all a1, a2 : A, eqA a1 a2 implies that eqB (f a1) (f a2).
Using our example, $\mathbb{N}/2$, f : $\mathbb{N}$to $\mathbb{N}$ is proper if inputs to f of the same parity produce outputs of the same parity. The successor function would be one example of such a proper function. When a term is composed of proper functions, it becomes possible to construct proofs for rewriting by equivalence relations in much the same way we can rewrite by equality. Coq provides automation in its standard library to support constructing these proofs. By using setoids in place of quotients, equivalence relations in place of equality, and proper functions in place of arbitrary functions, we mimic the functionality of quotients using setoids.
Our paper deals with repair across quotient type equivalences. To begin, we define type isomorphisms:
Definition 2.4.
Types A and B are isomorphic if there exist functions f : A $\to$ B and g : B $\to$ A such that
  • $\forall$ (a : A), g (f a) = a
  • $\forall$ (b : B), f (g b) = b
Type equivalences are type isomorphisms with an additional adjoint property, which can be derived from any isomorphism. A quotient type equivalence is simply an equivalence between two quotient types. When using setoids to represent quotients, we must instead define a notion of a setoid equivalence:
Definition 2.5.
@@ 2. Quotients and Setoids
Two setoids (A, eqA) and (B, eqB) are equivalent if there exist functions f : A $\to$ B and g : B $\to$ A satisfying the following properties:
@@
The function g is said to be inverse to f.
Two setoids are equivalent if the quotient types they represent would be isomorphic. One example of a nontrivial setoid equivalence between the setoids ($\mathbb{N}$, $\mathbb{N}$) and (Bool, =) is given by the function isEven : $\mathbb{N}$to Bool, sending even numbers to true and odd numbers to false. This function has inverse g : Bool $\to \mathbb{N}$ mapping true to 0 and false to 1. f is proper, since isEven respects parity, and g is automatically proper because Bool uses equality as its equivalence relation. The other conditions, isEven (g b) = b and $\mathbb{N}$ (g (isEven n)) n, are both satisfied, and thus this indeed forms a setoid equivalence.
@@ 3. Approach: Proof Term Transformation
To reiterate, our goal is to perform proof repair across quotient type equivalences. Because we are working in Coq, which does not have quotient types, we represent quotients using setoids. Thus, given two types A and B which may either be types using standard equality or setoids using some equivalence relation, we wish to turn terms defined over A into terms defined over B which no longer have any references to A.
To make the proof repair problem more concrete, we consider a specific example from the Pumpkin Pi paper. In Figure 1, we see two representations of the natural numbers. The first is a unary representation, while the second is a binary representation. Suppose that a user has begun proof development using the unary representation, implementing functions, such as addition, and theorems, such as that 0 is an identity for addition. Then, suppose that the user needs to switch to the binary representation later in development. Proof repair takes the functions and theorems defined over the unary natural numbers and turns them into functions defined over the binary natural numbers with the same behavior. These functions and theorems will make no reference to the unary natural numbers, so that the type can be deleted from the codebase entirely.
We will review how Pumpkin Pi repairs proofs across type equivalences by directly transforming proof terms across those equivalences (Section 3.1). Then, we extend Pumpkin Pi’s transformation to support setoid equivalences (Section 3.2).
@@
3.1. Pumpkin Pi’s Transformation
{grammar}
[…]
¡t¿ ::= ¡v¿   —   ¡s¿   —   $\Pi$ (¡v¿ : ¡t¿) . ¡t¿   —   $\lambda$ (¡v¿ : ¡t¿) . ¡t¿   —   ¡t¿ ¡t¿   —   Ind (¡v¿ : ¡t¿){¡t¿,…,¡t¿}
—   Constr (¡i¿, ¡t¿)   —   Elim(¡t¿, ¡t¿){¡t¿,…,¡t¿}
((DepConstr, DepElim), ($\iota$, $\eta$))
for types on both sides of the equivalence. [^2] While Pumpkin Pi has an $\eta$ term, meant to deal with $\eta$-expansions of constructors applied to eliminators, it appears to us that this term is not necessary. The example of nontrivial $\eta$ given in the Pumpkin Pi paper can be rewritten to instead have nontrivial $\iota$ and trivial $\eta$, and we are not aware of any cases where this cannot be done. Further, while the type of $\iota$ is always clearly defined, what $\eta$ should be is unclear for many examples. Thus, for the purposes of this paper, $\eta$ will always be trivial and we will ignore it. DepConstr and DepElim are, respectively, constructors and eliminators for each type, termed dependent constructors and dependent eliminators by Pumpkin Pi. The constructors must generate the elements of the type, and the eliminator must specify how to consume an element produced by the constructors. These constructors and eliminators must take the same shape, even if A and B themselves have different shapes.
For the type of unary naturals in Figure 1, we can choose depConstr for nat to be O and S, the constructors for nat. However, to repair to the binary naturals, we need to provide dependent constructors for N. These constructors must correspond across the equivalence, so we arrive at constructors with these types:
Definition depConstrNZero : N.
Definition depConstrNSuc : N -> N.
These dependent constructors do not share the type signatures of N’s constructors, but rather are user-defined functions corresponding to nat’s constructors. Likewise, we can choose the eliminator for nat, nat_rect, as the depElim for nat, and the dependent eliminator for N then takes the shape of nat’s eliminator:
Definition depElimN : $\forall$ (P : N -> Type),
(P depConstrNZero) ->
($\forall$ n : N, P n -> P (depConstrNSuc n)) ->
$\forall$ n : N, P n.
We give a concrete example of repairing addition in Figure 3.
Definition addNat (a b : nat) :=
depElimNat
(fun _ => nat -> nat)
(fun b => b)
(fun a IH b => depConstrNatSuc (IH b))
a
b.
Definition addN (a b : N) :=
depElimN
(fun _ => N -> N)
(fun b => b)
(fun a IH b => depConstrNSuc (IH b))
a
b.
Figure 3. Repairing the addition function from unary (left) to binary (right) natural numbers.
The fact that both dependent eliminators must have the same shape even when the underlying types do not is exactly why we need the remaining element of the configuration: $\iota$. This gives the $\iota$-reduction rules, which specify how to reduce an application of a dependent eliminator to a dependent constructor. When the shape is of the underlying type is the same as the shape of the configuration components, as is true for for nat, this will be definitional—the proof assistant will handle it automatically. However, if the inductive structure has changed, as it has for N, this will be a propositional equality.
Once we have defined the components of the configuration, we are ready to do repair. First, the functions we wish to repair are converted to explicitly refer to the configuration terms. Sometimes this happens via manual user annotation, and sometimes this happens via custom unification machinery inside of Pumpkin Pi that does this automatically for some classes of changes (see Section 4.2). Then, we follow the syntactic transformation outlined in Figure 4. We will discuss what it means for the repaired term to be correct in Section 6.
@@ 3.1. Pumpkin Pi’s Transformation
Figure 4. Transformation for repair across $A\simeq B$ with configuration
((DepConstr, DepElim), (Eta, Iota)), from previous work (Ringer, 2021). Our work adapts and extends this transformation.